OpenStudy (lena772):

Jill has been watching the squirrels that live in her neighborhood. The number of squirrels changes each week. n f(n) 1 8 2 12 3 18 4 27 Which function best shows the relationship between n and f(n)? f(n) = 12(1.5)n f(n) = 12(1.5)n-1 f(n) = 8(1.5)n f(n) = 8(1.5)n-1

OpenStudy (the_fizicx99):

Which do you think it is

OpenStudy (the_fizicx99):

f(n) = f(1) • r^n-1

OpenStudy (lena772):

8*r^8-1

OpenStudy (lena772):

what is r

OpenStudy (lena772):

I'm guessing r is 1.5 and the answer is D

OpenStudy (the_fizicx99):

Nope, 8*1.5^4-1 8*1.5^3 = ? It is D. Could you solve?

OpenStudy (the_fizicx99):

The exponent is time

OpenStudy (the_fizicx99):

Geddit? :D

OpenStudy (lena772):

yes thanks

OpenStudy (anonymous):

@tHe_FiZiCx99 witch was the answer

OpenStudy (the_fizicx99):

What?

OpenStudy (anonymous):

whats the answer

OpenStudy (anonymous):

Jill has been watching the squirrels that live in her neighborhood. The number of squirrels changes each week. n f(n) 1 8 2 12 3 18 4 27 Which function best shows the relationship between n and f(n)? f(n) = 12(1.5)n f(n) = 12(1.5)n-1 f(n) = 8(1.5)n f(n) = 8(1.5)n-1 was it D

OpenStudy (the_fizicx99):

-_- Lol really? You're actually asking me for the answer to a problem I explained? Look at the explanation.. and get you're answer through that. @Lena772 was able to understand it, I'm not giving you the answers..

OpenStudy (anonymous):

its D LOL

OpenStudy (the_fizicx99):

Then why'd you ask??

OpenStudy (anonymous):

I'm guessing r is 1.5 and the answer is D cause she said this and then you said Nope, 8*1.5^4-1 8*1.5^3 = ? It is D. Could you solve?

OpenStudy (the_fizicx99):

Yes, what didn't you get?

OpenStudy (anonymous):

Nope, 8*1.5^4-1 8*1.5^3 = ? It is D. Could you solve? you said nope

OpenStudy (the_fizicx99):

What she got wrong is that she thought it was 8-1.. ^n-1 is to find which ever value you want. Suppose you wanted to find for the 3rd week you would subtract 3-1 = 2 So, 8(1.5)^2 8*2.25 = 18

OpenStudy (anonymous):

oh lol okay can you help me in one more

OpenStudy (the_fizicx99):

Which?

OpenStudy (anonymous):

Zoe keeps track of the miles per gallon her car gets per week. She has accumulated the following data: (1, 24), (2, 24.38), (3, 24.76), (4, 25.14) What is the common difference or ratio? The common difference is 0.38. The common difference is 0.76. The common ratio is 0.99. The common ratio is 1.02.

OpenStudy (the_fizicx99):

This is arithmetic simply because their difference is .38 Subtract 24.76 - 24.38 = .38 Arithmetic is common difference. Geometric is common ratio.

OpenStudy (anonymous):

so its A

OpenStudy (the_fizicx99):

Yes..

OpenStudy (anonymous):

YAY LOL last one

OpenStudy (anonymous):

Melinda has shown that a function, f(x), increases by 4 for every unit in the domain. What does this prove? The function f(x) is an arithmetic sequence. The function f(x) is a geometric sequence. The function f(x) is not a sequence. This does not prove anything.

OpenStudy (the_fizicx99):

It increases by 4. is 4 geometric or arithmetic

OpenStudy (anonymous):

arithmetic

OpenStudy (the_fizicx99):

Ok

OpenStudy (anonymous):

so its A

OpenStudy (the_fizicx99):

Think of a graph..  have a better idea? Does that look geometric?